OpenStudy (anonymous):

Solve x + 2y = 13 4x + 8y = -9 (3, 5) (-1, 7) No Solutions Infinitely Many Solutions

OpenStudy (anonymous):

So i already used elimination and got... x + 2y = 13 + 4x + 8y = -9 __________________________ 5x + 10y = 4 but i don't know what to do next.

OpenStudy (anonymous):

No, that's not right, for elimination to work, first you pick a variable to ELIMINATE, to do that you take the 2 equations and multiply it by numbers that will give you opposites, so lets say you want to eliminate x, multiply the first equation by -4, and the second by 1, does that make sense?

OpenStudy (nincompoop):

how is that exactly an elimination method?

OpenStudy (nincompoop):

the more appropriate question would be, what did you eliminate in that method?

OpenStudy (anonymous):

@iambatman ok i understand. so -4(x+2y) = 13 do you multiplay 13 by -4 as well?

OpenStudy (anonymous):

Yes, the whole equation

OpenStudy (nincompoop):

his name is Bruce, btw

OpenStudy (anonymous):

Nin quit giving out my secret identity, you so phat.

OpenStudy (wolf1728):

Bruce Wayne, the billionaire financier that donates generously to the charities of Gotham City?

OpenStudy (anonymous):

so when both the equations cancel out then what? -4x - 8y = -52 4x + 8y = -9

OpenStudy (anonymous):

what is happening hahahahah

OpenStudy (anonymous):

@wolf1728 Yup, that's me ^.^ @kellyconnelly Yes, now add them just like the last problem so we can eliminate the variables, and look and x and y, you notice anything?

OpenStudy (anonymous):

at*

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

theyre zero?

OpenStudy (anonymous):

Yes, you get 0 = -61, meaning the answer is no solutions :)

OpenStudy (anonymous):

wow you're perfect

OpenStudy (anonymous):

thats exactly why

OpenStudy (wolf1728):

and to give a serious answer here, if you were to graph the 2 equations, they would NEVER intersect indicating no solution.

OpenStudy (anonymous):

Yes, they would be parallel lines, 

OpenStudy (wolf1728):

Hmmm, so does that mean the answer is No Solutions OR Infinitely Many Solutions ?

OpenStudy (anonymous):

No solutions, only way I could see it being infinitely many solutions is if they are coincident lines.

OpenStudy (wolf1728):

Okay, thanks. I just wanted to make absolutely sure.

OpenStudy (anonymous):

Np ^.^

OpenStudy (anonymous):

Just an fyi, it would be infinite solutions if you got something like 0 = 0.